Starting 3AM Thursday 29th October: Following on from our planned work in September (https://aastatus.net/37312) we are wanting to reboot some of our other core switches as we still have some oddities with them which is preventing is from bring back one of our BT links in to service. We will minimise the the disruption as much as possible but there may be short periods where broadband connections will drop and reconnect. We expect the work to be completed by 5:30AM and will update this post during the work.
This work has been rescheduled to happen on Thursday 29th October, 3AM.
This work has started, we have 3 switch reboots to do, and will be doing the first switch now.
Not a good start, the switch is failing to boot. We will halt this work and will be replacing the switch.
Staff are on their way to the datacentre with spare equipment. The switch being down is affecting a very small number of wholesale resellers (who have been contacted by email) but otherwise service is unaffected.
Replacement switch is configured and installed in to the rack and we're starting to bring it in to service. This may be a little disruptive as the network re-converges.
Replacing the switch did cause disruption as raised in https://aastatus.net/37984 - the switch is still powered down, and we'll reschedule this work again soon.
We have an understanding of the problems that occurred this morning and we have planned to bring the switch back in to service on Saturday 31st October between 2AM and 5AM. This work will be tracked on: https://aastatus.net/37985